OpenStudy (anonymous):

What is the solution to ? A. x = 5 or x = –7 B. x = –5 or x = 7 C. x = 5 or x = 7 D. no solutio

OpenStudy (anonymous):

equation is X^2-35=2x

OpenStudy (johnweldon1993):

We can subtract 2x from both sides to make the equation = 0 , then use the quadratic formula x² - 2x - 35 = 0 can you do this?

OpenStudy (anonymous):

\[x^2 -35 = 2x\] \[x^2-2x-35=0\] Find the roots You will end up with... \[(x-7)(x+5)=0\] So when is this equation true?

OpenStudy (anonymous):

okay @johnweldon1993 @DoYourHomework thank you

OpenStudy (johnweldon1993):

no problem!

OpenStudy (anonymous):

so x = 5 or x = 7?

OpenStudy (johnweldon1993):

not quite....remember those numbers have to multiply together to = -35 7 Times 5 doesn't equation -35 ...it does = 35 One of those numbers has to be a negative...Look again what @DoYourHomework wrote above :)

OpenStudy (anonymous):

x = 5 or x = –7 or x = –5 or x = 7